Western Family Steakhouse offers a variety of low-cost meals and quick service. Other than management, the steakhouse operates with two full-time employees who work 8 hours per day. The rest of the employees are part-time employees who are scheduled for 4-hour shifts during peak meal times. On Saturdays the steakhouse is open from 11:00 A.M. to 10:00 P.M. Management wants to develop a schedule for part-time employees that will minimize labor costs and still provide excellent customer service. The average wage rate for the part-time employees is $6.1 per hour, but the temp agency managing the part time staff will charge the steakhouse one extra dollar per hour for shifts starting after 3:00 PM. The total number of full-time and part-time employees needed varies with the time of day as shown.

Time Total Number
of Employees Needed
11:00 a.m. - Noon 7
Noon - 1:00 p.m. 8
1:00 p.m. - 2:00 p.m. 9
2:00 p.m. - 3:00 p.m. 6
3:00 p.m. - 4:00 p.m. 5
4:00 p.m. - 5:00 p.m. 5
5:00 p.m. - 6:00 p.m. 8
6:00 p.m. - 7:00 p.m. 11
7:00 p.m. - 8:00 p.m. 10
8:00 p.m. - 9:00 p.m. 8
9:00 p.m. - 10:00 p.m. 4

One full-time employee comes on duty at 11:00 A.M., works 4 hours, takes an hour off, and returns for another 4 hours. The other full-time employee comes to work at 1:00 P.M. and works the same 4-hours-on, 1-hour-off, 4-hours-on pattern.

  1. Develop a minimum-cost schedule for part-time employees. What is the total payroll for the part-time employees? If required, round your answer to the nearest dollar.

    Total daily salary cost = $  

    How many part-time shifts are needed?



    Use the surplus variables to comment on the desirability of scheduling at least some of the part-time employees for 3-hour shifts.




  2. Assume that part-time employees can be assigned either a 3-hour or a 4-hour shift. Develop a minimum-cost schedule for the part-time employees. If your answer is zero, enter "0".

    Optimal schedule for part-time employees:
    Starting Time Number of part-
    time employees
    (4-hour shifts)
    Number of part-
    time employees
    (3-hour shifts)
    11:00 A.M.
    12:00 P.M.
    1:00 P.M.
    2:00 P.M.
    3:00 P.M.
    4:00 P.M.
    5:00 P.M.
    6:00 P.M.
    7:00 P.M.

    How many part-time shifts are needed?



    What is the cost savings compared to the previous schedule? If required, round your answer to the nearest cent.

    Total cost reduced to $  .
Answer #1

Total Payroll For Part Time Employees

Time Total Number
of Employees Needed
Emplyees requirenent per hour Wages for Part Time(PT)
11:00 a.m. - Noon 7 1(full time) + 6 (part time) 6*6.1=$36.6
Noon - 1:00 p.m. 8 1(full time) + 7 (part time) 7*6.1=$42.7
1:00 p.m. - 2:00 p.m. 9 2(full time) + 7 (part time) 7*6.1=$42.7
2:00 p.m. - 3:00 p.m. 6 2(full time) + 4 (part time) 4*6.1=$24.4
3:00 p.m. - 4:00 p.m. 5 1(full time) + 4 (part time) 4*7.1=$28.4
4:00 p.m. - 5:00 p.m. 5 2(full time) + 3 (part time) 3*7.1=$21.3
5:00 p.m. - 6:00 p.m. 8 1(full time) + 7 (part time) 7*7.1=$49.7
6:00 p.m. - 7:00 p.m. 11 2(full time) + 9 (part time) 9*7.1=$63.9
7:00 p.m. - 8:00 p.m. 10 2(full time) + 8 (part time) 8*7.1=$56.8
8:00 p.m. - 9:00 p.m. 8 1(full time) + 7 (part time) 7*7.1=$49.7
9:00 p.m. - 10:00 p.m. 4 1(full time) + 3 (part time) 3*7.1=$21.3
TOTAL WAGES FOR PT $437.5